How do you find the slope and y intercept of an equation like this:
x=-2/3
Thanks!

Answer by jsmallt9(3759)

When working with lines and linear equations, you need to learn about a special group of lines. These lines have equations of the form:
x = some number
These equations are equations of vertical lines. Vertical lines are special because:
  • They have undefined slope. (Since slope = rise/run and vertical lines have no run (they do not go horizontally at all), the denominator of the slope is zero and dividing by zero is undefined.)
  • Only x = 0 has a y-intercept... and it is entire line that intersects/intercepts the y-axis be cause x=0 is the y-axis!. All the other "x = some number" equations have no y-intercepts
  • They cannot be written in slope-intercept form, y = mx+b

Because vertical lines and their equations, x = some number, are so exceptional and so different from the other lines and equations of lines, it is best just to learn the above about these lines and not try to use the "regular" forms and techniques on them.

Your equation, y = 2/3, is of this form. So it has undefined slope and, since it is not x = 0, it has no y-intercepts. It is a vertical line running through 2/3 on the x-axis.